# Fractional Programming for Stochastic Precoding over Generalized Fading Channels

### v1

@@ VI-B Test Results
Moreover, we compare the speed of convergence for the different algorithms. We now draw Algorithm 1 and Algorithm 2 separately. First, we evaluate the speed in iterations. As shown in Fig. 4, WMMSE cannot even yield monotonic improvement for the objective value. It can be seen that SWMMSE converges slowly; this low efficiency can be expected since online optimization mines for solution from massive channel samples. In particular, note that Algorithm 2 is even slower than Algorithm 1. Actually, this is not a surprising result. When we talk about acceleration in Section V, we mean the simplification of each iteration by eliminating matrix inverse. However, as shown in [40], this simplification is at the cost of slowing down the convergence in iterations. (As shown in [40], this drawback can be remedied by incorporating momentum into the FP method, but this is beyond the scope of this work.) Nevertheless, the acceleration in computation outweighs the deceleration in iterations. In plain words, even though it now takes more iterations to converge, the run time of each iteration becomes much shorter, and hence the overall time consumption is still reduced significantly by Algorithm 2. The results from Fig. 5 agree with the above discussion. Observe that Algorithm 2 is about 3 times faster than Algorithm 1. Actually, when Algorithm 1 just finishes one iteration, Algorithm 2 already reaches convergence. We further compare their run times in Table II by using different numbers of antennas. Observe that the efficiency advantage of Algorithm 2 over Algorithm 1 becomes larger when the number of antennas increases, so Algorithm 2 is well suited for the large-scale MIMO network.

### v2

@@ VI-B Test Results
Moreover, we compare the speed of convergence for the different algorithms. We now draw Algorithm 1 and Algorithm 2 separately. First, we evaluate the speed in iterations. As shown in Fig. 3 and Fig. 4, the convergence curves of the lower bounds are consistent with those of the actual objective function value. Observe also that WMMSE cannot even yield monotonic improvement for the objective value. It can be seen that SWMMSE converges slowly; this low efficiency can be expected since online optimization data-mines for solution from massive channel samples. In particular, note that Algorithm 2 is even slower than Algorithm 1. Actually, this is not a surprising result. When we talk about acceleration in Section V, we mean the simplification of each iteration by eliminating matrix inverse. However, as shown in [29], this simplification is at the cost of slowing down the convergence in iterations. Nevertheless, the acceleration in computation outweighs the deceleration in iterations. In plain words, even though it now takes more iterations to converge, the run time of each iteration becomes much shorter, and hence the overall time consumption is still reduced significantly by Algorithm 2. The results from Fig. 5 and Fig. 6 agree with the above discussion. Observe that Algorithm 2 is about 3 times faster than Algorithm 1. Actually, when Algorithm 1 just finishes one iteration, Algorithm 2 already reaches convergence. We further compare their run times in Table II by using different numbers of antennas. Observe that the efficiency advantage of Algorithm 2 over Algorithm 1 becomes larger when the number of antennas increases, so Algorithm 2 is well suited for the large-scale MIMO network.
@@
Furthermore, following [29], we accelerate Algorithm 2 by incorporating momentum into it. As shown in Fig. 7, momentum can indeed improve the convergence of Algorithm 2. Both Algorithm 1 and Algorithm 2 guarantee convergence to a stationary point of the lower bound problem. Since Algorithm 2 is based on a looser bound, its convergence slows down, but its per-iteration complexity is lower because of the cheaper subproblem. Thus, Algorithm. 2 is preferable when the computational power is highly limited and the full convergence is not required.
Lastly, we consider the case in which the rate weight varies from user to user. For simplicity, we focus on the single-cell network with the user moving speed $v=90$ km/h. The rate weights of the eight users are $\{0.617,0.094,0.171,0.887,0.886,0.251,0.102,0.772\}$. The simulation results are summarized in Fig. 8 and Fig. 9. It can be seen that the distinct-weight case does not differ much from the unit-weight case as formerly shown in Fig. 4 and Fig. 6. Thus, the proposed algorithm accounts for various rate weight settings.